OpenStudy (anonymous):

12. The temperature was 56 degrees this morning but dropped one degree per hour through the rest of the day. Use the graph you made for the equation for this problem, to approximate the answer to the following question: What temperature was it 15 hours later, according to your graph? 13. The temperature was 56 degrees this morning, but dropped one degree per hour through the rest of the day. Read over the problem and look at the graph you made to decide how to answer this question. Would it make sense to predict the temperature in 45 hours? Why or why not?

OpenStudy (anonymous):

12. 41degrees

OpenStudy (anonymous):

10. The temperature was 56 degrees this morning but dropped one degree per hour through the rest of the day. Write the linear equation for this problem in y=mx+b form. Note: this equation has a negative slope (m). Answer is: y= -x+56 11. The temperature was 56 degrees this morning but dropped one degree per hour through the rest of the day. Make a T-table for the equation for this problem, using the x-values 0, 5, 10. Type in the ordered pairs you get in the T-table as your answer. answer is: (0,56) (5,51) (10,46)

OpenStudy (anonymous):

No because it would drop more at night time due to the lack of sun Then possibly rise again the next day Also doesn't take into account weather change

OpenStudy (anonymous):

Eliza is correct for question number 13. Over 45 hours the temperature would have increased and decreased twice due to rise and fall of the sun, plus varying weather conditions. 45 hours is nearly 2 full days.

OpenStudy (anonymous):

For question 12: First find the slope, by taking two points from the relation and finding change in y over change in x. Like if one point is (a,b) and another is (x,y), the slope is (y - b) over (x - a). Then use whatever form you have for the equation of a line (point slope is easiest: if (a,b) is one of the points in the relation, and m is the slope you got, the equation is y - b = m(x - a)